Na2CO3(aq) + CuCl2(aq) =CuCO3(s) + 2NaCl(s) what is the theoretical yield of cuco3
1
Expert's answer
2019-10-28T07:49:44-0400

First of all, the equation is incorrect. NaCl will not be formed in a solid form as well as CuCO3 would, most probably, not be formed:


"2Na_2CO_3(aq)+2CuCl_2(aq)+H_2O \rightarrow (CuOH)_2CO_3(s)+CO_2(g)+4NaCl(aq)"


Now, let assume that the reaction still follows the initial equation. Thus, the theoretical yield is the relation between the amount in moles of the formed CuCO3 precipitate and the starting reagent consumed in the reaction (Na2CO3 or CuCl2):


"\gamma(CuCO_3)=\frac{n(CuCO_3)}{n(Na_2CO_3)}=\frac{n(CuCO_3)}{n(CuCl_2)}"


"\gamma(CuCO_3)=\frac{m(CuCO_3)M(Na_2CO_3)}{m(Na_2CO_3)M(CuCO_3)}=\frac{m(CuCO_3)M(CuCl_2)}{m(CuCl_2)M(CuCO_3)}"